OpenStudy (anonymous):

function if f'(x)=(f(x))^2 and f(0)=-1/5

OpenStudy (azureilai):

f(x)=x-1/5 f'(x) = (x-(1/5))(x-(1/5)) now you multiply them and combine like terms to get f'(x)

OpenStudy (anonymous):

thanks

OpenStudy (azureilai):

no prob